A person walks 46 m East and then walks 39 m at an angle of 47 degrees North of East. What is the magnitude of the total dis-placement ? Answer in units of m.
1
Expert's answer
2020-10-01T10:09:57-0400

We have two displacements: r1 (when a person walks 46 m East), r2 (when a person walks 39 m at an angle 47 degrees) and total displacement r.

Displacement along the x-axis:

"r_{1x} = 46\;m"

"r_{2x} = 39\; m\times cos(47º) = 39\; m\times0.68 = 26.52\; m"

"r_x = r_{1x} + r_{2x} = 46 \;m + 26.52\;m = 72.52\; m"

Displacement along the x-axis:

"r_{1y} = 0\; m"

"r_{2y} = 39\; m\times sin(47º) = 39\; m\times0.73 = 28.47 \;m"

"r_y = r_{1y} + r_{2y} = 0 + 28.47\;m = 28.47\; m"

Pythagorean theorem:

"r^2 = r^2_y + r^2_x"

"D = \sqrt{r^2_y + r^2_x} = \sqrt{72.52^2 + 28.47^2} = 77.90 \;m"

Answer: 77.90 m